ноги — н,can you produce a mechanism

for the top two
Answer #1

HC ca C- CH₂ + O2 = CH – ch3 H₂ C- C - CH₂ - CH₂ 17/H (= CH – CH₂ H₂C - (- CH₂ – CH3 = k In the first step the sp-Carbon atomHe- C = C- CH₂ H He -e = CH – CH₂ | - - -* TO + C-C Cu 3 CH HC - & - CH₂ - CH₂ In the first step is the formation of Carbocat